Answer with all steps and explaintion please:

♡ Evaluate. Does it diverge (В) е
Answer #1

o= simila) put 2- Sime when da= coso do whitiny = -1/2 & when n=1 = 0= 1/2 Above becomes T22 Si-simta Loso de - - 77 s costoHowever, we lan enpress interms of infinite senes 3 e?. 1+27 7 을 11 T 2! 3. n! hะ b * ex² It x2 + 24 +26, M8 2! 3! n=0 Now in